OpenStudy (anonymous):

Find dy/dx by using implicit differentiation. (xy)^(1/2) = 9+(x^2)(y)

OpenStudy (turingtest):

at what point are you stuck?

OpenStudy (anonymous):

using product rule on x^(1/2)* y ^(1/2) = 9 + (x^2)(y) (1/2)x^(-1/2) * y^(1/2) + (1/2)y^(-1/2)* dy/dx * x^(1/2) = x^2* dy/dx + y * 2x - now make dy/dx the subject - i've struggled a bit with that - maybe another opinion would be helpful

OpenStudy (anonymous):

 whew... not done.. i gotta rest my hand...